Status Update

An update on an existing published issue

Shared hosting issue

Date: 
Fri, 16/03/2012 - 09:18

We are experiencing an issue with our shared hosting platform this morning which has resulted in all shared hosting services being unavailable.

The file system on the server went into read-only mode around 09:15 this morning (we are still trying to ascertain why this happened). At the moment the server is performing a file system check to check the integrity of the data. Once this is completed then the server will be back online.

We are working hard to resolve this issue and hope to resume normal service soon. Please accept our apologies for any inconvenience caused.

Scheduled Maintenance: VPS Platform Upgrade

Date: 
Sat, 10/03/2012 - 23:00 - Sun, 11/03/2012 - 01:00

We will be upgrading the Citrix XenServer hypervisor that powers our VPS platform in this maintenance window. This upgrade will allow us to support new operating systems and provides a number of stability and performance enhancements.

Any disruption should be minimal but all VPSs will be at risk for the duration of this window.

Scheduled maintenance: Data Centre Migration

Date: 
Fri, 27/01/2012 - 21:00 - Sat, 28/01/2012 - 07:00

In conjunction with engineers from Gyron, we will be migrating all hardware at our primary site, Telstra’s London Hosting Centre, to Gyron’s own Centro Data Centre in Hemel Hempstead. All customer services will be affected during this window.

Sending email servers should queue your inbound email for delivery until the migration is complete. No email should be lost.

No IP addresses, hardware, software or DNS names will change during the migration - this is a physical relocation only - so you don’t need to update any configurations.

Scheduled Maintenance: VPN appliance software update

Date: 
Sun, 16/10/2011 - 22:00 - 22:30

We will be performing scheduled maintenance during this window to upgrade the software on our VPN remote access appliance. Only SSL VPN remote access and secure file transfer services will be affected, and not production web or email traffic.

We apologise for any inconcenience this may cause.

Update: This maintenance has been rescheduled.

Update 16/10/2011 22:41: This maintanence is still ongoing and we expect it to be completed shortly.

XenServer issue

Date: 
Sun, 27/03/2011 - 16:59

We have an issue with one of our XenServer hosts which is affecting some customer machines. We are working on the issue and customer virtual machines should be back online shortly.

We apologise for any inconvenience caused.

Update 17:15
All customer virtual machines are now booting and we are looking into the cause of the issue.

Update 23:15

RESOLVED: Support tickets when emailing support@networkeq.co.uk

Date: 
Sat, 05/03/2011 - 17:23

Due to the hardware issue we experienced with our IronPort, currently support tickets are not automatically generated when you email support@networkeq.co.uk. We will be manually making tickets for those customers emailing support@networkeq.co.uk in the meantime or you can raise a ticket on our website at https://secure.networkeq.co.uk/clients/submitticket.php?step=2&deptid=5

Update 22:00 08/03/2011
The IronPort appliance has now been replaced and this issue is now resolved.

Network issue

Date: 
Wed, 09/02/2011 - 16:42

We are currently experiencing an unexpected issue with one of our firewalls in Docklands, this has caused a loss of connectivity to all customer services. We are currently working on bringing the firewall back online and we will then investigate the root cause.

Please accept our apologies for any inconvenience.

Update 17:10: We have turned off the firewall in the cluster that we believe to be faulty and are currently working on bringing the other firewall back online. This issue is being dealt with at the highest priority, thank you for your patience.

Networking issue

Date: 
Sat, 05/02/2011 - 17:04

We are currently experiencing an issue with one switch in our virtual chassis which has caused our storage network to become unavailable, this is having a knock on effect on all services.

We are currently working to get the switch back online and resume normal service. We will post updates here as we know more.

Please accept our apologies for any inconvenience.

XenServer host issue

Date: 
Wed, 19/01/2011 - 16:59

We are currently experiencing an unexpected issue with one of our XenServer hosts. This is affecting a number of client virtual machines. We are currently investigating the problem and, where possible, moving client machines to other hosts.

We will post updates here when we have more information. Please accept our apologies for any inconvenience.

Hosted Exchange Issue

Date: 
Tue, 04/01/2011 - 13:30 - 14:05

We are an aware of an issue affecting connectivity to our Hosted Exchange environment.

We are looking into the issue and will provide an update shortly. We apologise for any invoncenience caused.

Update 14:05: We believe this issue is now resolved and will provide further details shortly.

Status Update Feed